# Benchmarking Mental State Representations in Language Models

### v1

@@ 4.4 Contrastive Activation Addition
We finally compare models’ accuracy on three BigToM tasks in Table 1. Each model has been evaluated three times: without any intervention, using ITI, and using CAA. Hyperparameter details can be found in Section A.6. Note that we use steering vectors computed using the Forward Belief task for all three tasks to test their generalisability (Zhu et al., 2024). As can be seen from the table, performance without intervention is generally lower across tasks and model sizes, with the larger Llama-2-70B and Llama-2-70B-chat models exhibiting higher accuracy. Performance for Pythia models of different sizes does not change much, with the fine-tuned Pythia-6.9B-chat showing better performance on single true belief (TB) and false belief (FB) tasks but not on their conjunction (Both). ITI demonstrates modest improvements over no intervention for Llama-2 models. Improvements for Pythia models are consistent and higher, up to $+17$. The only exception is Pythia-6.9B-chat, for which ITI is not always beneficial.
@@
CAA consistently delivers the most substantial accuracy improvements across all models and tasks, up to $+56$ for Llama-2-13B-chat on the (Backward Belief), which Gandhi et al. have identified as the hardest task. Despite its relatively small size, Llama-2-13B-chat excels in all three tasks when using CAA. Larger 70B models often achieve accuracies close to or exceeding $90\%$. Smaller models like Pythia-70M and Pythia-410M also show significant gains with CAA, though the absolute performance is still lower than Llama-2. Overall, our results indicate that it is possible to effectively enhance ToM reasoning in LMs without needing to train any probe, which yields even improved results. Furthermore, we show that CAA steering vectors generalise well, yielding substantial performance gains across all ToM tasks.
5 Limitations and Future Work
@@
Our study focused on expanding experiments from the model perspective, examining architectures, sizes, fine-tuning, and prompt design, all within the same dataset. A natural extension of our work is replicating these experiments across multiple datasets and more model families. Given the rapid pace of new language model releases, studying all available models is impractical, particularly considering computational resource constraints. Nevertheless, our approach can be adopted to support new benchmarks or to evaluate newly released models as they become available.
6 Conclusion
Our study addresses a significant gap in understanding LMs by investigating their internal representation of mental states. We conducted an extensive benchmark involving various LM types, sizes, fine-tuning approaches, and prompt designs to examine the robustness of these representations. Our findings reveal that scaling LMs’ size and, in particular for smaller LMs, fine-tuning are key to developing representations of others’ beliefs. We are the first to demonstrate that such prompt variations influence model representations, and we also demonstrate the feasibility of enhancing models’ ToM reasoning by steering their activations without training any probe. Overall, our work contributes valuable insights into the factors influencing LMs’ mental state representations, shedding light on avenues for improving their performance in ToM tasks.
Societal Impact
While our work is foundational and remains distant from specific applications with direct societal impact, it’s important to recognise the ethical implications of modelling and predicting mental states. Handling sensitive aspects of individuals’ inner experiences and emotions requires careful consideration to avoid reinforcing biases or misunderstanding psychological nuances.
Author Contributions Statement
M. Bortoletto designed and directed the project, performed the experiments, analysed the results, and wrote the manuscript.
C. Ruhdorfer provided feedback on the experimental results and contributed to editing the final manuscript.
L. Shi contributed to proofreading the final manuscript.
A. Bulling supervised the project and contributed to the writing of the manuscript.
@@ Acknowledgements
L. Shi was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C. Ruhdorfer. The authors would like to especially thank Hsiu-Yu Yang and Anna Penzkofer for their support and the numerous insightful discussions.
We used PyTorch (Paszke et al., 2019) to manipulate tensors, Transformers (Wolf et al., 2020) to evaluate the models and NNsight (Fiotto-Kaufman, ) to perform activation interventions. We performed our data analysis using NumPy (Harris et al., 2020), Pandas (pandas development team, 2020; Wes McKinney, 2010), and SciPy (Virtanen et al., 2020). We made our plots using Matplotlib (Hunter, 2007) and seaborn (Waskom, 2021).
@@ A.1.1 BigToM
BigToM (Gandhi et al., 2023) is constructed using GPT-4 (Achiam et al., 2023) to populate causal templates and combine elements from these templates. Each causal template is set up with a context and a description of the protagonist (e.g. “Noor is working as a barista […]”), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”). The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”). The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs – true or false. Gandhi et al. (2023) generated 200 templates and extracted 25 conditions from each template, resulting in 5,000 test samples. In this work, following Zhu et al. (2024) and Gandhi et al. (2023), we focused on the 6 most important conditions, corresponding to true and false beliefs on the following three tasks:
@@
The dataset was released under the MIT license and can be accessed at https://github.com/cicl-stanford/procedural-evals-tom. We report one example for each task in the boxes below, where the text defining true belief or false belief task is shown in blue and red, respectively.
Example of Forward Belief Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk. Noor does not see her coworker swapping the milk.
Does Noor believe the milk pitcher contains oat milk or almond milk?
a) Noor believes the milk pitcher contains oat milk.
b) Noor believes the milk pitcher contains almond milk.
Example of Forward Action Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor sees her coworker swapping the milk. Noor does not see her coworker swapping the milk.
What will Noor do?
a) Noor will make the cappuccino using the milk in the pitcher.
Example of Backward Belief Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor opens the fridge again and reaches for the oat milk. Noor makes the cappuccino using the milk in the pitcher.
Does Noor believe the milk pitcher contains oat milk or almond milk?
a) Noor believes the milk pitcher contains oat milk.
b) Noor believes the milk pitcher contains almond milk.
@@ A.1.2 Linear Probes
Our probing approach is illustrated in Figure 5. We cache activations both at the attention and residual stream level. In our experiments, we use residual stream activations. To perform ITI and compare it to CAA, we also cache attention heads’ activations. We trained the probes using the L-BFGS solver (Liu & Nocedal, 1989) with L2 penalty with inverse of regularisation strength $10$ for a maximum of $1000$ iterations. We use zero as random seed.
@@ A.1.3 Language models
Table 2: The 12 models used in this work. The checkmark indicates we additionally study the fine-tuned (Chat) version of the model.
LM Size + Chat Tokens $d_{model}$ Layers
Llama-2 (Touvron et al., 2023) 7B 2T 4096 32
@@ A.1.4 Examples of Prompt Variations
Default prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: Noor believes the milk pitcher contains oat milk.
Random prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Misleading prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: Noor believes the milk pitcher contains oat milk.
Belief: The water valve is closed.
Time specification prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: In the end, Noor believes the milk pitcher contains oat milk.
Initial belief prompt Story: Noor is working as a barista at a busy coffee shop. Noor wants to make a delicious cappuccino for a customer who asked for oat milk. Noor grabs a milk pitcher and fills it with oat milk. Noor believes that the milk pitcher contains oatmilk. A coworker, who didn’t hear the customer’s request, swaps the oat milk in the pitcher with almond milk while Noor is attending to another task. Noor does not see her coworker swapping the milk.
Belief: Noor believes the milk pitcher contains oat milk.
@@ A.5 Inference-time intervention
Inference-time intervention (Li et al., 2023c, ITI) employs a two-step process. First, it trains a probe for each attention head across all layers of a LM. These probes are evaluated on a validation set, and the top-$k$ heads with the highest accuracy are selected. Subsequently, during inference, ITI steers the activations of these top heads along the directions defined by their corresponding probes. Formally, ITI can be defined as an additional term to the multi-head attention:
$x_{l+1}=x_{l}+\sum_{h=1}^{H}Q_{l}^{h}\left(\mathrm{Att}_{l}^{h}(P_{l}^{h}x_{l}){\color[rgb]{0.5,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,1}\pgfsys@color@cmyk@stroke{0.50}{1}{0}{0}\pgfsys@color@cmyk@fill{0.50}{1}{0}{0}\ +\ \alpha\sigma_{l}^{h}\theta_{l}^{h}}\right)$
@@
A.6 Activation editing hyperparameters
Table 3 reports results obtained on the three BigToM tasks with the hyperparameters used for ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023). We report an example of prompt used for evaluation in the box below.
Prompt used for downstream evaluation (Section 4.4) Answer the questions based on the context. Keep your answer concise, few words are enough, maximum one sentence. Answer as ’Answer:<option>)<answer>’.
Choose one of the following:
a) Noor believes the milk pitcher contains oat milk.
b) Noor believes the milk pitcher contains almond milk.
Answer:
Table 3: Comparison between ITI (Li et al., 2023c) and CAA (Rimsky et al., 2023) on the three BigToM tasks (Gandhi et al., 2023). For ITI, the subscript indicates the value of the coefficient used. For CAA, the subscript indicates first the value of the coefficient used and second the layer at which intervention takes place.
@@ A.6 Activation editing hyperparameters
TB FB TB $\land$ FB TB FB TB $\land$ FB TB FB TB $\land$ FB

### v3

@@ 4 Results
We compare models’ accuracy on three BigToM tasks in Table 1 (Llama) and Table 3 (Pythia). Each model has been evaluated three times: without any intervention, using ITI, and using CAA. Hyperparameter details can be found in Section A.6. Note that we use steering vectors computed using the Forward Belief task for all three tasks to test their generalisability.
Performance without intervention is generally lower across tasks and model sizes, with the larger Llama-2-70B and Llama-2-70B-chat models exhibiting higher accuracy. Performance for Pythia models of different sizes does not change much, with the fine-tuned Pythia-6.9B-chat often showing better performance on single true belief (TB) and false belief (FB) tasks but not on their conjunction (Both).
ITI demonstrates modest improvements over no intervention for Llama-2 models. Improvements for Pythia models are consistent and higher, up to $+17$. The only exception is Pythia-6.9B-chat, for which ITI is not always beneficial.
@@
CAA consistently delivers the most substantial accuracy improvements across all models and tasks, up to $+56$ for Llama-2-13B-chat on the Backward Belief task, which Gandhi et al. have identified as the hardest task. Despite its relatively small size, Llama-2-13B-chat excels in all three tasks when using CAA. Larger 70B models often achieve accuracies close to or exceeding $90\%$. Smaller models like Pythia-70M and Pythia-410M also show significant gains with CAA, though the absolute performance is still lower than Llama-2. To further demonstrate CAA’s effectiveness, we applied it while evaluating models on a control task where the causal event in the story is replaced by a random one that does not change the environment (e.g., A musician starts playing music while Noor is making the latte). Table 4 shows improved results for all models, indicating that CAA improves performance on ToM tasks without compromising the models’ ability on control tasks.
Overall, our results indicate that it is possible to further enhance ToM reasoning in LMs in a computationally cheap way, without needing to train any probe. Furthermore, we show that the CAA steering vectors are general, yielding substantial performance gains across all ToM tasks.
5 Discussion and Conclusion
In this work, we conducted extensive experiments across 12 LMs to examine their internal representation of beliefs of self (oracle) and others (protagonist). Our experiments show similar emergence patterns across all the models we evaluated (RQ1): oracle belief representations generally form in the first layers, while for protagonist they emerge at the intermediate layers. Moreover, probing accuracy increases with model size and, more crucially for smaller models, with fine-tuning (RQ1) (Figure 3). While larger models show higher probing accuracy, this could be due to their higher dimensionality – at the same time increasing the number of learning parameters in the probes and offering more spurious patterns to fit. To control for this, we ran two experiments: one using randomly permuted labels, and one projecting activations onto their top-$k$ principal components to reduce probe size. Results show that high-dimensional probes cannot learn random label mappings (Fig. 8), and that reduced representations retain most of the original accuracy (Fig. 4, 11, 12). Together, these findings suggest that probes capture structured belief representations rather than spurious correlations (RQ2). We then explore if these representations are robust to prompt variations. Our experiments demonstrate that LMs possess robust belief representations when taking an omniscient perspective (Fig. 10), whereas their representations of others’ beliefs are more brittle (RQ3), with probing accuracy decreasing for semantically neutral prompts (Fig. 5, 9). Our final set of experiments shows that belief representations can be strengthened using CAA (RQ4). CAA steers model activations in a generalisable way, significantly improving performance across multiple ToM tasks while being computationally cheaper than ITI (Table 1, 3). For instance, with Llama-2-70B, ITI requires training 5,120 probes (64 attention heads $\times$ 80 layers), whereas CAA only needs 80 vectors, one per layer.
In summary, our key takeaway is that while models can robustly represent beliefs from an omniscient perspective,
representations of others’ beliefs improve with model size and fine-tuning, are structured yet brittle – but also easily steerable.
Together, our findings suggest several promising directions for future work. Better understanding the similar emergence pattern of belief representations across LMs can inform architecture design and training strategies. Especially for smaller models, future work could explore how different types of fine-tuning (e.g., human feedback vs. synthetic data) influence the emergence of internal belief representations. Demonstrating that these representations are structured rather than spurious validates the use of probing as a meaningful tool to study how LMs’ represent beliefs of self and others, and encourages internal model analysis as part of evaluation pipelines. However, the brittleness of belief representations to prompts – particularly when attributing beliefs to others – suggests that the perspective-taking machinery needed for robust ToM reasoning remains fragile, and highlights the need for robustness benchmarks and new approaches to improve generalisation. Finally, our success with CAA shows that belief representations can be strengthened in a generalisable and efficient way, opening up opportunities for real-time model steering in socially grounded tasks. While CAA offers a post-hoc remedy, future research should also explore methods for directly embedding perspective-taking circuits into model architectures.
Limitations
@@ 5 Discussion and Conclusion
Our study focused on expanding experiments from the model perspective, examining architectures, sizes, fine-tuning, and prompt design, all within the same dataset. A natural extension of our work is replicating these experiments across multiple datasets and more model families. Given the rapid pace of new language model releases, studying all available models is impractical, particularly considering computational resource constraints. Nevertheless, our approach can be adopted to support new benchmarks or to evaluate newly released models as they become available. Finally, while in this work we focused on beliefs, our experimental approach can be adapted to investigate how LMs represent desires, emotions, intentions, or preferences. Future research exploring other types of mental states can use our findings to determine whether similar or distinct patterns emerge.
@@ Acknowledgements
L. Shi was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C. Ruhdorfer.
@@ A.1.1 BigToM
BigToM (Gandhi et al., 2023) is constructed using GPT-4 (Achiam et al., 2023) to populate causal templates and combine elements from these templates. Each causal template is set up with a context and a description of the protagonist (e.g. “Noor is working as a barista […]”), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”). The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”). The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs – true or false. Gandhi et al. (2023) generated 200 templates and extracted 25 conditions from each template, resulting in 5,000 test samples. In this work, following Zhu et al. (2024) and Gandhi et al. (2023) we focused on the 6 most important conditions, corresponding to true and false beliefs on the following three tasks:
@@
The dataset was released under the MIT license and can be accessed at https://github.com/cicl-stanford/procedural-evals-tom. We report one example for each task in Example A.9, A.9, and A.9, where the text defining true belief or false belief task is shown in blue and red, respectively.
@@ A.1.2 Linear probes
Our probing approach is illustrated in Figure 6. For our experiments, we cache activations at the residual stream level. To perform ITI and compare it to CAA, we also cache attention heads activations. We trained the probes using the L-BFGS solver (Liu and Nocedal, 1989) with L2 penalty with inverse of regularisation strength $10$ for a maximum of $1000$ iterations. We use zero as random seed.
@@ A.1.3 Language models
LM Size + SFT + RLHF Tokens $d_{model}$ Layers
Llama-2 7B 2T 4096 32
@@ A.1.4 Examples of prompt variations
Example A.9 shows an example of Original prompt. Examples of prompt variations are provided in Example A.9 (Random), Example A.9 (Misleading), Example A.9 (Time Specification), and Example A.9 (Initial Belief).
@@ A.2 Model size and fine-tuning
A.2.1 Overfitting Issues
Figure 3 also that probing accuracy at early layers is particularly low across all models, performing even worse than random. This happens due to overfitting, which may be caused by spurious features introduced by the initial coding strategy of language models, where individual token representations are mixed together (Gurnee et al., 2023). We also identified the same issue when reproducing the results in Zhu et al. (2024), who address it by manually clipping all accuracies below random chance to 50%. [^4] https://github.com/Walter0807/RepBelief/blob/0fc86396f2f0a998643ea01786eb3db4dd20ff9c/probe.py#L60 Since probing experiments require training a large number of probes for each model, both we and Zhu et al. (2024) trained each probe for the same fixed number of epochs (1,000). However, for activations from the earlier layers, overfitting occurs very quickly - often within the first 10 iterations.
We ran an experiment with Llama2-7B-chat, reducing training to fewer than 10 iterations, and found that the probes performed at random chance. Therefore, to fully resolve this issue, we would need to choose the number of training epochs for each probe individually. This would likely flatten the observed "U" shape in the results. However, this process would be computationally expensive and does not contribute to our main research questions. Rather than artificially adjusting accuracies to 50%, we prefer to present the results as they are.
@@ A.2.1 Overfitting Issues
(a) Original activations.
(b) $k=1000$ largest principal components of the activations.
(c) $k=100$ largest principal components of the activations.
Figure 8: Comparison between accuracy on belief probing and accuracy obtained on a control task.
@@ A.3 Sensitivity to prompting
Accuracy on protagonist belief probing for Pythia models is shown in Figure 9.
@@ A.4 Dimensionality reduction
Probing accuracy obtained by Pythia models for the protagonist setting is reported in Figure 11.
@@
Figure 11: We compare the probing accuracy obtained by using the original set of activations (All) with the accuracy obtained by considering only the first $n=\{2,10,100,1000\}$ principal components. For Pythia: All(70m) = $512$, All(410m) = $1024$, All(1b) = $2048$, All(6.9b) = $4096$, All(12b) = $5120$. Results for oracle are shown in Figure 12.
@@ A.5 Inference-time intervention
Inference-time intervention (Li et al., 2023b, ITI) employs a two-step process. First, it trains a probe for each attention head across all layers of a LM. These probes are evaluated on a validation set, and the top-$k$ heads with the highest accuracy are selected. Subsequently, during inference, ITI steers the activations of these top heads along the directions defined by their corresponding probes. Formally, ITI can be defined as an additional term to the multi-head attention:
$x_{l+1}=x_{l}+\sum_{h=1}^{H}Q_{l}^{h}\left(\mathrm{Att}_{l}^{h}(P_{l}^{h}x_{l}){\color[rgb]{0.56,0.27,0.52}\definecolor[named]{pgfstrokecolor}{rgb}{0.56,0.27,0.52}\ +\ \alpha\sigma_{l}^{h}\theta_{l}^{h}}\right)$
@@
A.6 Activation editing
Table 3 reports results obtained on the three BigToM tasks with the corresponding hyperparameters used for ITI (Li et al., 2023b) and CAA (Rimsky et al., 2023). We report an example of prompt used for evaluation in Example A.9. Table 4 shows the accuracy obtained by using CAA on the Forward Belief True Control task in BigToM. On this control task, CAA produced improved results for all model, proving that CAA not only improves performance on ToM tasks, but also does not degrades the models’ ability to perform other tasks.
@@ A.6 Activation editing
TB FB Both TB FB Both TB FB Both
@@
Table 3: Activation intervention: comparison between ITI (Li et al., 2023b) and CAA (Rimsky et al., 2023). For ITI, the subscript indicates the value of the coefficient $\alpha_{\mathrm{ITI}}$ used: $\mathrm{Acc}_{\alpha_{\mathrm{ITI}}}$. For CAA, the subscript indicates first the value of the coefficient $\alpha$ used and second the layer $l$ at which intervention takes place: $\mathrm{Acc}_{\alpha_{\mathrm{CAA}},l}$.
Model Method Control CAA Parameters
Llama-2-7b No int. $44$
CAA $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+22}}$ $2.0$, $11$
Llama-2-7b-chat No int. $56$
CAA $70_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+14}}$ $1.0$, $11$
Llama-2-13b No int. $52$
CAA $85_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+33}}$ $2.0$, $12$
Llama-2-13b-chat No int. $84$
CAA $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}$ $1.0$, $12$
Llama-2-70b No int. $90$
CAA $99_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$ $2.0$, $16$
Llama-2-70b-chat No int. $69$
CAA $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+23}}$ $1.5$, $18$
Pythia-70m No int. $41$
CAA $62_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+21}}$ $1.0$, $2$
Pythia-410m No int. $48$
CAA $67_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+19}}$ $2.0$, $4$
Pythia-1b No int. $44$
CAA $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+15}}$ $2.0$, $8$
Pythia-6.9b No int. $44$
CAA $56_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $1.5$, $12$
Pythia-6.9b-chat No int. $55$
CAA $68_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}$ $1.5$, $15$
Pythia-12b No int. $44$
CAA $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $2.0$, 00
Table 4: Results for CAA Rimsky et al. (2023) on the Forward Belief True Control task in BigToM Gandhi et al. (2023). Numbers indicate accuracy scores, with differences (CAA -- No int.) as subscripts.